OpenStudy (vishweshshrimali5):

If e^y + xy = e then find the value of y'' at y = 1

OpenStudy (vishweshshrimali5):

y'' means second derivative of y w.r.t. x

OpenStudy (anonymous):

Clarification needed: Is \(e\) approximately \(2.71\) or a variable?

OpenStudy (vishweshshrimali5):

Well sorry , I don't have any idea regarding that. but the answer says 1/e^2

OpenStudy (vishweshshrimali5):

hope that this may help u

OpenStudy (anonymous):

I unfortunately cannot see how to solve this. Someone other than me needs to give it a try; it seems like a differential equation.

OpenStudy (anonymous):

Maybe implicit differentiation?

OpenStudy (anonymous):

Hmm, I still have an x in the equation at the end..

OpenStudy (anonymous):

\[\frac{d}{dx}[e^y+xy=e] \rightarrow y'e^y+xy'+y=0.\]

OpenStudy (anonymous):

\[\frac{d}{dx}[y'e^y+xy'+y=0] \rightarrow y'y'e^y+y''e^y+xy''+y'+y'=0.\] \[y'=\frac{-y}{(e^y+x)}\] Can anyone check my work up to this point? I think from here, it's \[y''=\frac{-2y'-(y')^2e^y}{e^y+x}\] then substitute in y' and y..

OpenStudy (vishweshshrimali5):

y' = ?

OpenStudy (anonymous):

Oh, duh! I forgot to simply solve for x. x=0 when y =1, so yeah all that's left is y''=1/e^2.